Consider the following 2-person economy. Output per day Water balloons Art projects Ben 15 3 Evelyn 12 2 a. For Ben and Evelyn, what is the opportunity cost of an art project? Who has a comparative advantage? (Hint: Opportunity Cost A = Production B/Production A) b. Who has the comparative advantage in the production of water balloons? c. Suppose that both Ben and Evelyn decide to work 1 day, and each specializes in their area of competitive advantage. How many of each good will be produced?
Answer #1

a) Opportunity cost of an art project for

Ben = 15/3= 5

Evelyn = 12/2 = 6

Ben has a comparative advantage in the production of Art project

b) Opportunity cost of the water balloons for

Ben = 3/15 = 0.2

Evelyn = 2/12 = 0.16

Evelyn has a comparative advantage in the production of water balloons

c) Ben will produce art project to a total of 3

Evelyn will produce water balloons to a total of 12